# Non-Hermitian Linear Electro-Optic Effect Through Interactions of Free and Bound Charges

### v1

@@ V.2 Indefinite gain response
In order to explain the directional dependence of the gain response of the material, next we characterize the tensor $\mathbf{\overline{\varepsilon}}^{\prime\prime}\equiv\frac{\mathbf{\overline{\varepsilon}}-\mathbf{\overline{\varepsilon}}^{\dagger}}{2i}$, which determines the power exchanged between the material and the wave. Lannebère et al. (2022, 2025b, 2025a). For the model of Eq. (33), it is given by $\mathbf{\overline{\varepsilon}}^{\prime\prime}=\mathrm{Im}\left(\varepsilon_{d}\right)\mathbf{1}_{3\times 3}+\mathbf{\overline{\chi}}_{\text{NH}}$ where
@@
For a standard passive material, all the eigenvalues of $\mathbf{\overline{\varepsilon}}^{\prime\prime}$ are required to be positive to guarantee that the energy of the wave is irreversibly dissipated in the material. Interestingly, in our problem $\mathbf{\overline{\chi}}_{\text{NH}}$ is an indefinite matrix, meaning that it has both positive and negative eigenvalues. Specifically, its nontrivial eigenvalues are ±$\pm\frac{1}{2}|\alpha_{0}|$, with $\alpha_{0}=\mathrm{Re}(\varepsilon_{c})-3i\mathrm{Im}(\varepsilon_{c})$, and the corresponding eigenvectors are ±$\pm\frac{\alpha_{0}}{|\alpha_{0}|}\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$. The positive eigenvalue corresponds to an electro-optic response that results in increased dissipation, whereas the negative eigenvalue to an electro-optic response that results in reduced dissipation.
As $\mathbf{\overline{\chi}}_{\text{NH}}$ trivially commutes with the response of the unbiased material (a scalar), the eigenvalues of $\mathbf{\overline{\varepsilon}}^{\prime\prime}$ are $\lambda_{\pm}\equiv\mathrm{Im}(\varepsilon_{d})\pm\frac{1}{2}|\alpha_{0}|$, corresponding to the eigenvectors ±$\pm\frac{\alpha_{0}}{|\alpha_{0}|}\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$, and $\lambda_{0}\equiv\mathrm{Im}(\varepsilon_{d})$ corresponding to the eigenvector $\hat{\mathbf{e}}_{y}$. Thus, in the presence of an electric bias the level of dissipation depends on the wave polarization in the material. In our system, the parameter $\alpha_{0}$ is predominantly real-valued. Thereby, the polarizations that activate the maximum (minimum) dissipation are such that $\mathbf{E}\sim s\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$ and $\mathbf{E}\sim-s\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$, with $s={\mathop{\rm sgn}}\left({{\mathop{\rm Re}\nolimits}\left\{{{\varepsilon_{c}}}\right\}}\right)$. This explains why the maximum dissipation/gain occurs for angles $\theta=\mp\pi/4$: for such directions of propagation the electric field of the TM wave is roughly oriented along ±$\pm\pi/4$, matching closely one of the eigenvectors of $\mathbf{\overline{\varepsilon}}^{\prime\prime}$.
Eventually, when the bias velocity is strong enough such that in some frequency range $\frac{1}{2}|\alpha_{0}|>\mathrm{Im}(\varepsilon_{d})$, the smallest eigenvalue $\lambda_{-}$ can become negative. In such a case, the material exhibits an indefinite gain response Lannebère et al. (2025b), as it can provide both dissipative and gain interactions, depending on the wave polarization. When $\lambda_{-}<0$, the electromagnetic waves polarized as $-\frac{\alpha_{0}}{|\alpha_{0}|}\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$ are amplified.
@@
Figure 4 depicts $\lambda_{-}$ as a function of frequency for different values of the bias velocity. In the unbiased case (green dashed curve), $\lambda_{-}=\mathrm{Im}(\varepsilon_{d})$ is positive for all frequencies, consistent with the passivity of the material. When the bias velocity is increased beyond some threshold, $\lambda_{-}$ can reach negative values, indicating that the associated eigenpolarization can be amplified by the material. As seen, the electro-optic response is stronger in the low-frequency band, where $\lambda_{-}$ is more negative than in the high-frequency band. In particular, for the case corresponding to the stability threshold of the closed system (red dotted curve), $\lambda_{-}$ only reaches negative values in the low-frequency band whereas it is positive everywhere in the high-frequency band. It is interesting to note that the system can exhibit gain ($\lambda_{-}<0$) even in the regime where the closed system is stable ($\omega^{\prime\prime}\leq 0$ for all real-valued wave vectors). This occurs because the eigenpolarization of the TM-polarized plane wave, $\mathbf{E}_{\text{TM}}$, is not close enough to $-\frac{\alpha_{0}}{|\alpha_{0}|}\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$, the eigenvector associated with $\lambda_{-}$ that triggers the material gain. The transverse optical phonon frequency $\omega_{0}$ of natural materials is typically in the range of the tens of THz Kittel (2004). Thus, the NHEO effect should be the strongest for frequencies in the THz range.
@@ V.4 Faraday-type rotation and electromagnetic isolation
To this end, we derived the transmission matrix $\mathbf{\overline{T}}$ that relates the transverse components of the incident $\mathbf{E}_{t}^{\text{inc}}=\left(E_{x}^{\text{inc}}\;E_{z}^{\text{inc}}\right)^{T}$ and transmitted $\mathbf{E}_{t}^{\text{tr}}=\left(E_{x}^{\text{tr}}\;E_{z}^{\text{tr}}\right)^{T}$ electric fields as $\mathbf{E}_{t}^{\text{tr}}=\mathbf{\overline{T}}\cdot\mathbf{E}_{t}^{\text{inc}}$. It is shown in appendix C that for normal incidence (incident wave propagates along $+y$) the transmission matrix for a material slab (infinitely extended along $x$ and $z$) of thickness $d$ is
@@
where $\beta_{j}=\left[\cos(k_{j}d)-i\frac{(\varepsilon_{j}+\varepsilon_{\text{ext}})}{2\sqrt{\varepsilon_{j}}\sqrt{\varepsilon_{\text{ext}}}}\sin(k_{j}d)\right]^{-1}$ with $j=+$ or --, $\varepsilon_{\text{ext}}$ the permittivity of the surrounding dielectric (e.g., air) and the remaining parameters defined as in Eq. (37).
@@ V.5 Power beating
Next, we examine the Poynting vector for a superposition of plane waves. The electric field for a superposition of $+$ and -- modes propagating along the $+y$ direction is given by
@@
${\bf{S}}=\frac{1}{{2\omega{\mu_{0}}}}{\mathop{\rm Re}\nolimits}\left\{{k_{+}^{*}{{\left|{{A_{+}}}\right|}^{2}}{e^{-2{k^{\prime\prime}_{+}}y}}+k_{-}^{*}{{\left|{{A_{-}}}\right|}^{2}}{e^{-2{k^{\prime\prime}_{-}}y}}-\frac{1}{3}\left({k_{+}+k_{-}^{*}}\right){A_{+}}A_{-}^{*}{e^{i\left({{k_{+}}-k_{-}^{*}}\right)y}}}\right\}\hat{\mathbf{e}}_{y}.$ (40)
We write the wave numbers in terms of real and imaginary parts: $k_{\pm}=k^{\prime}_{\pm}+ik^{\prime\prime}_{\pm}$. As seen, similar to what happens in other non-Hermitian systems Lannebère et al. (2022); Buddhiraju et al. (2020); Hadad and Engheta (2020); Fernandes et al. (2024), the Poynting vector is the sum of the power transported by the $+$ and -- waves, plus an interference term that oscillates with spatial frequency $k^{\prime}_{+}-k^{\prime}_{-}$.
@@
Figure 8: Normalized Poynting vector as a function of the propagation distance $y$ for (a) $\omega\approx 0.224\omega_{p}$ at the edge of the low-frequency band associated with the -- mode. (b) $\omega\approx 1.364\omega_{p}$ at the edge of the high-frequency band associated with the $+$ mode. The parameters of the simulation are $A_{+}=A_{-}=1$, $\Gamma=\gamma\to 0^{+}$ and the remaining parameters are as in Fig. 3. The electric field polarizations associated with the points where the curve slope either vanishes or is a maximum/minimum are represented in (a). The field polarizations in case (b) are nearly identical.
The two operating frequencies of Fig. 8 are coincident with the cutoff frequencies of the -- ($+$) mode in the low (high)-frequency band, respectively (see Fig. 6). Such a choice ensures that $|k_{+}-k_{-}|$ is maximized in each band and provides the highest rate of gain/loss per unit of length. As seen, similar to an ideal MOSFET-metamaterial Lannebère et al. (2022), our system alternates between active and dissipative states depending on the wave polarization. The wave amplification and attenuation are a result of the energy exchanged between the dynamic fields and the static current. This exchange is rooted in nonlinear mechanisms, akin to the operation of a field-effect transistor Fernandes et al. (2024).
@@ VI Conclusion
In conclusion, this paper introduces a novel mechanism to implement the NHEO effect, achieving optical gain through nonlinear interactions between free and bound electrons. This mechanism operates independently of the anomalous velocity of Bloch electrons and the Berry curvature dipole. Our semiclassical phenomenological model shows that under non-equilibrium conditions and within specific low-symmetry systems, the linearized electromagnetic response becomes not only nonreciprocal but also capable of exhibiting polarization-dependent gain.
Our model ensures global passivity and adheres to the principles of microscopic reversibility. By considering the specific spatial symmetries of the model, we can significantly narrow down the permissible forms of nonlinear coupling. A notable aspect of the NHEO response is that it necessitates broken inversion symmetry. We demonstrate that an electrically biased material with 2mm group symmetries can exhibit a robust nonreciprocal and non-Hermitian response. We identify (natural) materials belonging to the class of polar/ferroelectric metals with the 2mm group symmetry, such as for example WTe2, as potential candidates to implement this effect experimentally.
@@ Appendix A Linearization of nonlinear conservative systems described by a Lagrangian
[…]
@@
The $N$ equations of motion [$N\equiv\text{dim}(\mathbf{q})$] of the system are obtained from the Euler-Lagrange equations Landau and Lifshitz (1980)
@@
where $\delta q_{j}$, $\delta\dot{q}_{j}$ and $\delta(\partial_{x}q_{j})$ are respectively the $j$th component of the vectors $\mathbf{q}-\mathbf{q}_{0}$, $\dot{\mathbf{q}}-\dot{\mathbf{q}}_{0}$ and $\partial_{x}{\mathbf{q}}-\partial_{x}{\mathbf{q}}_{0}$. The subscript 00 indicates that the functions are evaluated at the bias point. Substituting the Taylor expansions of $\frac{\partial\mathcal{L}_{\text{NL}}}{\partial q_{i}}$, $\frac{\partial\mathcal{L}_{\text{NL}}}{\partial\dot{q}_{i}}$ and $\frac{\partial\mathcal{L}_{\text{NL}}}{\partial(\partial_{x}q_{i})}$ in Eq. (42) and taking into account that the equilibrium state satisfies the Euler-Lagrange equations, $\left.\frac{\partial\mathcal{L}_{\text{NL}}}{\partial q_{i}}\right|_{0}=\frac{\partial}{\partial t}\left(\left.\frac{\partial\mathcal{L}_{\text{NL}}}{\partial\dot{q}_{i}}\right|_{0}\right)+\frac{\partial}{\partial x}\left(\left.\frac{\partial\mathcal{L}_{\text{NL}}}{\partial(\partial_{x}q_{i})}\right|_{0}\right)$, we find that the linearized equations of motion are
@@ Appendix B Generalized microscopic model
$m_{b}\left(\partial_{tt}+\gamma\partial_{t}+\sum_{j=x,y,z}\omega_{0j}^{2}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\right)\mathbf{P}$ $=q^{2}n_{b}\left(\mathbf{E}+\mathbf{\overline{C}}_{21}\cdot qn\mathbf{v}\right)$ (49c)
Here, $n$ is the free electron density, which is related to the velocity through the continuity equation [Eq. (49b)]. It can be easily checked that provided $\mathbf{\overline{C}}_{12}^{T}=-\mathbf{\overline{C}}_{21}$ [Eq. (18)], the fields are still constrained by a conservation law of the type $\nabla\cdot{{\bf{S}}_{{\rm{tot}}}}+{\partial_{t}}{W_{{\rm{tot}}}}=-Q$, with
${{\bf{S}}_{{\rm{tot}}}}$ $={\bf{E}}\times{\bf{H}}+\frac{1}{2}nm\left({{\bf{v}}\cdot{\bf{v}}}\right){\bf{v}},$ (50a)
${W_{{\rm{tot}}}}$ $={W_{{\rm{EM}}}}+\frac{{{1}}}{{2}}nm{\bf{v}}\cdot{\bf{v}}+\frac{{{m_{b}}}}{{2{n_{b}}{q^{2}}}}\left({\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}+\sum\limits_{j}{\omega_{0j}^{2}{\bf{P}}\cdot{{{\bf{\hat{e}}}}_{j}}\otimes{{{\bf{\hat{e}}}}_{j}}}\cdot{\bf{P}}\right),$ (50b)
$Q$ $={{mn}\Gamma}{\bf{v}}\cdot{\bf{v}}+\gamma\frac{{{m_{b}}}}{{{n_{b}}{q^{2}}}}{\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}.$ (50c)
Hence, this model also ensures that the material response remains passive ($Q>0$). Note that now both the Poynting vector ($\mathbf{S}_{\rm{tot}}$) and the energy density ($W_{\rm{tot}}$) have field and matter components.

### v2

@@ V.2 Indefinite gain response
In order to explain the directional dependence of the gain response of the material, next we characterize the tensor $\mathbf{\overline{\varepsilon}}^{\prime\prime}\equiv\frac{\mathbf{\overline{\varepsilon}}-\mathbf{\overline{\varepsilon}}^{\dagger}}{2i}$, which determines the power exchanged between the material and the wave Lannebère et al. (2022, 2025b, 2025a). For the model of Eq. (33), it is given by $\mathbf{\overline{\varepsilon}}^{\prime\prime}=\mathrm{Im}\left(\varepsilon_{d}\right)\mathbf{1}_{3\times 3}+\mathbf{\overline{\chi}}_{\text{NH}}$ where
@@
For a standard passive material, all the eigenvalues of $\mathbf{\overline{\varepsilon}}^{\prime\prime}$ are required to be positive to guarantee that the energy of the wave is irreversibly dissipated in the material. Interestingly, in our problem $\mathbf{\overline{\chi}}_{\text{NH}}$ is an indefinite matrix, meaning that it has both positive and negative eigenvalues. Specifically, its nontrivial eigenvalues are $\pm\frac{1}{2}|\alpha_{0}|$, with $\alpha_{0}=\mathrm{Re}(\varepsilon_{c})-3i\mathrm{Im}(\varepsilon_{c})$, and the corresponding eigenvectors are $\pm\frac{\alpha_{0}}{|\alpha_{0}|}\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$. The positive eigenvalue corresponds to an electro-optic response that results in increased dissipation, whereas the negative eigenvalue to an electro-optic response that results in reduced dissipation.
As $\mathbf{\overline{\chi}}_{\text{NH}}$ trivially commutes with the response of the unbiased material (a scalar), the eigenvalues of $\mathbf{\overline{\varepsilon}}^{\prime\prime}$ are $\lambda_{\pm}\equiv\mathrm{Im}(\varepsilon_{d})\pm\frac{1}{2}|\alpha_{0}|$, corresponding to the eigenvectors $\pm\frac{\alpha_{0}}{|\alpha_{0}|}\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$, and $\lambda_{0}\equiv\mathrm{Im}(\varepsilon_{d})$ corresponding to the eigenvector $\hat{\mathbf{e}}_{y}$. Thus, in the presence of an electric bias the level of dissipation depends on the wave polarization in the material. In our system, the parameter $\alpha_{0}$ is predominantly real-valued. Thereby, the polarizations that activate the maximum (minimum) dissipation are such that $\mathbf{E}\sim s\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$ and $\mathbf{E}\sim-s\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$, with $s=\text{sgn}\left(\text{Re}\left\{\varepsilon_{c}\right\}\right)$. This explains why the maximum dissipation/gain occurs for angles $\theta=\mp\pi/4$: for such directions of propagation the electric field of the TM wave is roughly oriented along $\pm\pi/4$, matching closely one of the eigenvectors of $\mathbf{\overline{\varepsilon}}^{\prime\prime}$.
Eventually, when the bias velocity is strong enough such that in some frequency range $\frac{1}{2}|\alpha_{0}|>\mathrm{Im}(\varepsilon_{d})$, the smallest eigenvalue $\lambda_{-}$ can become negative. In such a case, the material exhibits an indefinite gain response Lannebère et al. (2025b), as it can provide both dissipative and gain interactions, depending on the wave polarization. When $\lambda_{-}<0$, the electromagnetic waves polarized as $\frac{\alpha_{0}}{|\alpha_{0}|}\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$ are amplified.
@@
Figure 4 depicts $\lambda_{-}$ as a function of frequency for different values of the bias velocity. In the unbiased case (green dashed curve), $\lambda_{-}=\mathrm{Im}(\varepsilon_{d})$ is positive for all frequencies, consistent with the passivity of the material. When the bias velocity is increased beyond some threshold, $\lambda_{-}$ can reach negative values, indicating that the associated eigenpolarization can be amplified by the material. As seen, the electro-optic response is stronger in the low-frequency band, where $\lambda_{-}$ is more negative than in the high-frequency band. In particular, for the case corresponding to the stability threshold of the closed system (red dotted curve), $\lambda_{-}$ only reaches negative values in the low-frequency band whereas it is positive everywhere in the high-frequency band. It is interesting to note that the system can exhibit gain ($\lambda_{-}<0$) even in the regime where the closed system is stable ($\omega^{\prime\prime}\leq 0$ for all real-valued wave vectors). This occurs because the eigenpolarization of the TM-polarized plane wave, $\mathbf{E}_{\text{TM}}$, is not close enough to $\frac{\alpha_{0}}{|\alpha_{0}|}\hat{\mathbf{e}}_{x}+\hat{\mathbf{e}}_{z}$, the eigenvector associated with $\lambda_{-}$ that triggers the material gain. The transverse optical phonon frequency $\omega_{0}$ of natural materials is typically in the range of the tens of THz Kittel (2004). Thus, the NHEO effect should be the strongest for frequencies in the THz range.
@@ V.4 Faraday-type rotation and electromagnetic isolation
To this end, we derived the transmission matrix $\mathbf{\overline{T}}$ that relates the transverse components of the incident $\mathbf{E}_{t}^{\text{inc}}=\left(E_{x}^{\text{inc}}\;E_{z}^{\text{inc}}\right)^{T}$ and transmitted $\mathbf{E}_{t}^{\text{tr}}=\left(E_{x}^{\text{tr}}\;E_{z}^{\text{tr}}\right)^{T}$ electric fields as $\mathbf{E}_{t}^{\text{tr}}=\mathbf{\overline{T}}\cdot\mathbf{E}_{t}^{\text{inc}}$. It is shown in appendix C that for normal incidence (incident wave propagates along +y+y) the transmission matrix for a material slab (infinitely extended along $x$ and $z$) of thickness $d$ is
@@
where $\beta_{j}=\left[\cos(k_{j}d)-i\frac{(\varepsilon_{j}+\varepsilon_{\text{ext}})}{2\sqrt{\varepsilon_{j}}\sqrt{\varepsilon_{\text{ext}}}}\sin(k_{j}d)\right]^{-1}$ with $j=+$ or -, $\varepsilon_{\text{ext}}$ the permittivity of the surrounding dielectric (e.g., air) and the remaining parameters defined as in Eq. (37).
@@ V.5 Power beating
Next, we examine the Poynting vector for a superposition of plane waves. The electric field for a superposition of ++ and - modes propagating along the +y+y direction is given by
@@
$\mathbf{S}=\frac{1}{2\omega\mu_{0}}\text{ Re}\left\{{k_{+}^{*}{{\left|{{A_{+}}}\right|}^{2}}{e^{-2{k^{\prime\prime}_{+}}y}}+k_{-}^{*}{{\left|{{A_{-}}}\right|}^{2}}{e^{-2{k^{\prime\prime}_{-}}y}}-\frac{1}{3}\left({k_{+}+k_{-}^{*}}\right){A_{+}}A_{-}^{*}{e^{i\left({{k_{+}}-k_{-}^{*}}\right)y}}}\right\}\hat{\mathbf{e}}_{y}.$ (40)
We write the wave numbers in terms of real and imaginary parts: $k_{\pm}=k^{\prime}_{\pm}+ik^{\prime\prime}_{\pm}$. As seen, similar to what happens in other non-Hermitian systems Lannebère et al. (2022); Buddhiraju et al. (2020); Hadad and Engheta (2020); Fernandes et al. (2024), the Poynting vector is the sum of the power transported by the ++ and - waves, plus an interference term that oscillates with spatial frequency $k^{\prime}_{+}-k^{\prime}_{-}$.
@@
Figure 8: Normalized Poynting vector as a function of the propagation distance $y$ for (a) $\omega\approx 0.224\omega_{p}$ at the edge of the low-frequency band associated with the - mode. (b) $\omega\approx 1.364\omega_{p}$ at the edge of the high-frequency band associated with the ++ mode. The parameters of the simulation are $A_{+}=A_{-}=1$, $\Gamma=\gamma\to 0^{+}$ and the remaining parameters are as in Fig. 3. The electric field polarizations associated with the points where the curve slope either vanishes or is a maximum/minimum are represented in (a). The field polarizations in case (b) are nearly identical.
The two operating frequencies of Fig. 8 are coincident with the cutoff frequencies of the - (++) mode in the low (high)-frequency band, respectively (see Fig. 6). Such a choice ensures that $|k_{+}-k_{-}|$ is maximized in each band and provides the highest rate of gain/loss per unit of length. As seen, similar to an ideal MOSFET-metamaterial Lannebère et al. (2022), our system alternates between active and dissipative states depending on the wave polarization. The wave amplification and attenuation are a result of the energy exchanged between the dynamic fields and the static current. This exchange is rooted in nonlinear mechanisms, akin to the operation of a field-effect transistor Fernandes et al. (2024).
@@ VI Conclusion
In conclusion, this paper introduces a novel mechanism to implement the NHEO effect, achieving optical gain through nonlinear interactions between free and bound electrons. This mechanism operates independently of the anomalous velocity of Bloch electrons and the Berry curvature dipole. Our phenomenological model describes both the nonlinearities arising from the anomalous velocity, as well as nonlinearities arising from the interband effects, which are not captured by the usual Boltzmann approximation. The model shows that in low-symmetry conductors under non-equilibrium conditions, the linearized electromagnetic response becomes not only nonreciprocal but also capable of exhibiting polarization-dependent gain, even in the absence of the anomalous velocity contribution.
Our model ensures global passivity and adheres to the principles of microscopic reversibility. By considering the specific spatial symmetries of the model, we can significantly narrow down the permissible forms of nonlinear coupling. A notable aspect of the NHEO response is that it necessitates broken inversion symmetry. We demonstrate that an electrically biased material with the symmetries of the 2mm point group can exhibit a robust nonreciprocal and non-Hermitian response. We identify (natural) materials belonging to the class of polar/ferroelectric metals with 2mm point group symmetry, such as for example WTe2, as potential candidates to implement this effect experimentally.
@@ Appendix A Linearization of nonlinear conservative systems described by a Lagrangian
$S=\int_{t_{1}}^{t_{2}}\int\mathcal{L}_{\text{NL}}\left(\mathbf{q},\dot{\mathbf{q}},\partial_{x}{\mathbf{q}}\right)dxdt$ (41)
@@
The $N$ equations of motion $N\equiv\text{dim}(\mathbf{q})$] of the system are obtained from the Euler-Lagrange equations Landau and Lifshitz (1980)
@@
where $\delta q_{j}$, $\delta\dot{q}_{j}$ and $\delta(\partial_{x}q_{j})$ are respectively the $j$th component of the vectors $\mathbf{q}-\mathbf{q}_{0}$, $\dot{\mathbf{q}}-\dot{\mathbf{q}}_{0}$ and $\partial_{x}{\mathbf{q}}-\partial_{x}{\mathbf{q}}_{0}$. The subscript 0 indicates that the functions are evaluated at the bias point. Substituting the Taylor expansions of $\frac{\partial\mathcal{L}_{\text{NL}}}{\partial q_{i}}$, $\frac{\partial\mathcal{L}_{\text{NL}}}{\partial\dot{q}_{i}}$ and $\frac{\partial\mathcal{L}_{\text{NL}}}{\partial(\partial_{x}q_{i})}$ in Eq. (42) and taking into account that the equilibrium state satisfies the Euler-Lagrange equations, $\left.\frac{\partial\mathcal{L}_{\text{NL}}}{\partial q_{i}}\right|_{0}=\frac{\partial}{\partial t}\left(\left.\frac{\partial\mathcal{L}_{\text{NL}}}{\partial\dot{q}_{i}}\right|_{0}\right)+\frac{\partial}{\partial x}\left(\left.\frac{\partial\mathcal{L}_{\text{NL}}}{\partial(\partial_{x}q_{i})}\right|_{0}\right)$, we find that the linearized equations of motion are
@@ Appendix B Generalized microscopic model
$m_{\text{b}}\left(\partial_{tt}+\gamma\partial_{t}+\sum_{j=x,y,z}\omega_{0j}^{2}\hat{\mathbf{e}}_{j}\otimes\hat{\mathbf{e}}_{j}\right)\mathbf{P}$ $=q^{2}n_{\text{b}}\left(\mathbf{E}+\mathbf{\overline{C}}_{21}\cdot qn\mathbf{v}\right)$ (49c)
Here, $n$ is the free electron density, which is related to the velocity through the continuity equation [Eq. (49b)]. It can be easily checked that provided $\mathbf{\overline{C}}_{12}^{T}=-\mathbf{\overline{C}}_{21}$ [Eq. (18)], the fields are still constrained by a conservation law of the type $\nabla\cdot{{\bf{S}}_{{\text{tot}}}}+{\partial_{t}}{W_{{\text{tot}}}}=-Q$, with
${{\bf{S}}_{{\text{tot}}}}$ $={\bf{E}}\times{\bf{H}}+\frac{1}{2}nm\left({{\bf{v}}\cdot{\bf{v}}}\right){\bf{v}},$ (50a)
${W_{{\text{tot}}}}$ $={W_{{\text{EM}}}}+\frac{{{1}}}{{2}}nm{\bf{v}}\cdot{\bf{v}}+\frac{{{m_{\text{b}}}}}{{2{n_{\text{b}}}{q^{2}}}}\left({\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}+\sum\limits_{j}{\omega_{0j}^{2}{\bf{P}}\cdot{{{\bf{\hat{e}}}}_{j}}\otimes{{{\bf{\hat{e}}}}_{j}}}\cdot{\bf{P}}\right),$ (50b)
$Q$ $={{mn}\Gamma}{\bf{v}}\cdot{\bf{v}}+\gamma\frac{{{m_{\text{b}}}}}{{{n_{\text{b}}}{q^{2}}}}{\partial_{t}}{\bf{P}}\cdot{\partial_{t}}{\bf{P}}.$ (50c)
Hence, this model also ensures that the material response remains passive ($Q>0$). Note that now both the Poynting vector ($\mathbf{S}_{\text{tot}}$) and the energy density ($W_{\text{tot}}$) have field and matter components.